OpenStudy (anonymous):

solve for a 10ac-x/11=-3

OpenStudy (anonymous):

10ac-x/11=-3 -10ac -10ac get a on one side - seperate from x -x/11 = -3 -10ac +3 +3 add 3 to both sides -x/11 +3 = -10ac divide everything by -10c to get a alone x/110c +3 = a

OpenStudy (anonymous):

Thank you, but I was trying to write 10ac-x as a numerator with 11 as a denominator that equals -3. I'm so confused! Thanks for your reply though.

OpenStudy (anonymous):

No problem sorry I couldn't be much help.

OpenStudy (anonymous):

Thats ok! Thanks anyway

OpenStudy (anonymous):

No problem :)